OpenStudy (anonymous):

Whats the missing number? –3.4 + (–8 + 2) = [ ? + (–8)] + 2

OpenStudy (freckles):

hint: a+(b+c)=(a+b)+c this is the associative property of addition

OpenStudy (freckles):

if you notice that property we are adding the same numbers up but just in a different order

OpenStudy (freckles):

like for a+(b+c) we are adding b and c first then we are going to add the a but for the other side (a+b)+c we are adding a and b first then we are adding the c

OpenStudy (freckles):

example of the associative property: (3+(-5))+7=3+(-5+7)

OpenStudy (anonymous):

@onedirection_ @mathsciencehistory @tylermcmullen23

OpenStudy (freckles):

so what do you think is the missing number?

OpenStudy (freckles):

–3.4 + (–8 + 2) = [ ? + (–8)] + 2 compare both sides what number is on the left that isn't on the right?

OpenStudy (anonymous):

freckles is practically giving u the answer

OpenStudy (anonymous):

its -3.4

OpenStudy (anonymous):

I'm in 6th grade i'm not sure what your trying to tell me

OpenStudy (anonymous):

lol look at it its in the same format on both sides of the equal sign

OpenStudy (freckles):

well I hope he will learn how to compare both sides of an equation now since he didn't really anything this time

OpenStudy (anonymous):

haha im sure thats girl

OpenStudy (anonymous):

but yea

OpenStudy (freckles):

I was just trying to help him compare both sides but since the answer is given now he doesn't really need to do anything

OpenStudy (anonymous):

lol its fine i dont think she wanted help